OpenStudy (anonymous):

what is equivalent to (4) / (3+ radical 2)? can you explain how you did it?

OpenStudy (gorv):

\[\frac{ 4 }{ 3+ \sqrt 2 }\]

OpenStudy (anonymous):

gorv

OpenStudy (gorv):

\[\frac{ 4 }{ 3+\sqrt2}*\frac{ 3-\sqrt2 }{ 3-\sqrt2 }\]

OpenStudy (anonymous):

Round each number to the nearest whole number. What is the best estimate for the sum of 317.38 82.21? A. 410 B. 400 C. 399 D. 390 go to my question

OpenStudy (gorv):

lol @b877vf[-0;odc u r spamming others Q ...... if u want help from anyone just righ @ followed by the name of the person

OpenStudy (anonymous):

Thank you gorv i understand it now.

OpenStudy (gorv):

welcome:)